OpenStudy (anonymous):

A force F applied to an object of mass m produces as acceleration of 3 m/s^2. The same force applied to a second object of mass M produces an acceleration of 1 m/s^2. (a) What is the ratio of m/M? (b) If m and M are combined, find their acceleration under the action of the force F.

OpenStudy (anonymous):

Attempt to the solution: a.) m/M = A/a= 1/3= 0.333 b.) F= m(3m/s^2) m= F/3m/s^2 F= M(1m/s^2 M= F/(1m/s^2)) m+M = F/3 m/s^2 + F/(1m/s^2) =F+3F/3m^2/s^4 m+M = 4F/(3m/s^2) F= ma m(3m/s^2)= [(4F)/(3m/s^2)](a) a= {[(F/3m/s^2)](3m/s^2)}/{(4F)/(3m/s^2)} a= 3 m/s^2/4 a= (3/4) m/s^2 is that correct?

OpenStudy (anonymous):

@sauravshakya :)

OpenStudy (anonymous):